OpenStudy (anonymous):

tell whether the sequence is arithmetic, geometric, or neither. find the next three terms of the sequence. 1/3, 0, 1, -2

OpenStudy (anonymous):

This one is surely not arithmetic because: 1/3 - 0 = 1/3 0 - 1 = -1 1 - 2 = -1 If the results were equal, it was going to be arithmetic. Here is a webpage that may help: http://www.purplemath.com/modules/series3.htm I'll continue working to check if it's geometric or not.

OpenStudy (anonymous):

It's not geometric too. Now we have to find the sequence pattern in order to find the next three terms. Notice that some instructors as to find the "previous" three terms.

OpenStudy (anonymous):

I want to help more.. but need to go to bed; it's too late here =$ Good luck ^^"